OpenStudy (anonymous):

x.y = 20 x.x.y.y = 328 x>y how : x.x.x-y.y.y ? x.x.x.x+y.y.y.y.y ?

OpenStudy (anonymous):

please , answer my question !!!!!!!!!!!!!

OpenStudy (anonymous):

your question is wrong. are you sure you typed it correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you say x.y = 20, you mean x times y =20 right?

OpenStudy (anonymous):

is the second eqn what u have given is correct??

OpenStudy (unklerhaukus):

\[xy = 20\]\[y=\frac {20} x\] \[x^2y^2 = 328\]\[x^2\left(\frac {20} x\right)^2=328\]\[20^2=328\]\[400=328???\]FALSE

OpenStudy (anonymous):

xy=20 x=20/y

OpenStudy (anonymous):

check your eqns.....

OpenStudy (anonymous):

wrong question

OpenStudy (unklerhaukus):

there are no solutions to the first two equations